OpenStudy (anonymous):

4/11z^2-11z - 9/11z-11 subtract and simplify

OpenStudy (anonymous):

4/11z^2-130/11z-11

OpenStudy (anonymous):

z(4/11z^3-130/11)-11

jhonyy9 (jhonyy9):

4/11z*2 -11z -9/11z-11 =4/z(11z-11) -9/11z-11 =4-9z/z(11z-11)